# On Geometric Evolution and Microlocal Regularity of the Navier–Stokes Equations

### v2

@@ IV Geometric Evolution Operators and Microlocal Representation of Solutions
  • $\mathcal{T}$ is the time-ordering operator.
  • $G(t,s)$ is the associated Green’s operator.
  • $\tilde{\phi}_{0}$ is the initial microlocal data.
@@
Finally, the physical solution is recovered by integrating over the cotangent directions
$\phi(x,t)=\int_{S_{x}^{\star}M}\tilde{\phi}(x,t,\xi)\,d\xi$ (78)
@@
which defines a geometrically canonical reconstruction of the velocity field. The reconstruction equation (78) recovers $\phi$ in a weak or averaged sense, which is sufficient for the purposes of the present analysis.
This formulation allows for the study of the system’s regularity through the microlocal behavior of $\tilde{\phi}$, reducing the global analysis of Navier-Stokes to the dynamical properties of the flow induced on $S^{*}M$. Analogously, we follow the same procedure for vorticity, obtaining the following equation
$\partial_{t}\tilde{\omega}+\mathcal{L}_{X_{u}}\tilde{\omega}=-\nu\|\xi\|_{g^{\prime}}^{2}\tilde{\omega}+\tilde{R}_{\omega}+i\xi\wedge\tilde{F},\qquad\xi\wedge\tilde{\omega}=0$ (79)
@@
The condition $\xi\wedge\tilde{\omega}=0$ is the microlocal version of $d\omega=0$. Equation (79) is linear and dissipative, and its solution is formally written as
$\tilde{\omega}(t)=\mathcal{T}\bigg[\exp\left[-\int_{t_{0}}^{t}(\mathcal{L}_{X_{u}}+\nu\|\xi\|_{g^{\prime}}^{2})\,ds\right]\tilde{\omega}_{0}\bigg]\\ +\int_{t_{0}}^{t}G(t,s)\big[\tilde{R}\omega(s)+i\xi\wedge\tilde{F}(s)\big]ds$ (80)
If $\tilde{\phi}(t)$ is a solution to (77), then $\tilde{\omega}(t):=d\tilde{\phi}(t)$ satisfies the microlocal vorticity equation (79). This is because $d(dP)=0$ and $X_{u}$ depends on $u$ and its derivatives, the exterior derivative does not commute exactly with the Lie derivative along $X_{u}$, a direct computation yields the commutator identity
$d,\mathcal{L}_{X_{u}}]=\mathcal{L}_{dX_{u}}$ (81)
When acting on microlocal scalar amplitudes $\tilde{\phi}$, the right–hand side produces lower–order terms depending on $\nabla u$, and we obtain the estimate
$d,\mathcal{L}_{X_{u}}]\tilde{\phi}=\mathcal{O}(\nabla u)\tilde{\phi}$ (82)
The viscous contribution in the lifted equation is modeled by the multiplication operator
$D:=\nu\|\xi\|_{g^{\prime}}^{2}$ (83)
whose commutator with $d$ is given explicitly by
$d,D]=d(\nu\|\xi\|_{g^{\prime}}^{2})\wedge(\cdot)$ (84)
Combining the transport and dissipative contributions, we define the total commutator operator
$\left[d,\mathcal{L}_{X_{u}}+\nu\|\xi\|_{g^{\prime}}^{2}\right]={R}$ (85)
which is a zeroth–order operator controlled by the geometry and the velocity field. More precisely, there exists a constant $C>0$ such that
$\|{R}_{u}\|_{L^{2}\to L^{2}}\leq C\left(\|\nabla u\|_{L^{\infty}}+1\right)$ (86)
As a consequence, if $\tilde{\phi}$ is a solution of the lifted scalar equation, then the microlocal vorticity $\tilde{\omega}:=d\tilde{\phi}$ satisfies the corresponding lifted equation up to lower–order controlled terms, which can be absorbed into the energy estimates. And applying the exterior derivative to (69) yields exactly (79) with a solution reconstructed also from
$\omega(x,t)=\int_{S_{x}^{*}M}\tilde{\omega}(x,t,\xi)\,d\xi$ (87)
It is inferred that the microlocal lift of the Navier-Stokes equations produces a well-defined dissipative dynamical system on the cosphere bundle, whose solution admits an explicit representation through evolution operators. The exact compatibility between the velocity and vorticity equations guaranties the consistency of the formalism.
@@
This approach reduces the global regularity problem to the control and convergence of geometric microlocal invariants, the analysis of which is addressed in the following section.
V Stability, Energy and Geometric Criteria for Regularity and Smoothness
@@
In the previous sections, the incompressible Navier–Stokes equations were reformulated as a linear, non-autonomous, dissipative dynamical system acting on microlocal lifts of differential forms over the compact phase space $S^{*}M$. We now exploit this structure to derive quantitative stability estimates and geometric criteria for regularity [28]. The central idea of this section is that any potential singular behavior of the physical solution must manifest itself as a geometric instability of the microlocal dynamics on $S^{*}M$ [29, 7]. Viscosity, compactness, and the geometric nature of the lift combine to impose strong restrictions on such instabilities.
@@
For those reasons, throughout this section, $(M,g)$ denotes a smooth Riemannian manifold and $S^{*}M$ its unit cosphere bundle endowed with the canonical Liouville measure $\mu_{S^{*}M}$. We assume that the velocity field satisfies
$u\in L^{\infty}([t_{0},t];W^{2,\infty}(M))$ (88)
@@
and that all geometric quantities associated with the effective connection $\nabla^{u}$ remain smooth on $t_{0},t]\times S^{*}M$. Let $\lambda\in\{\phi,\omega\}$ denote either the velocity $1$-form or the vorticity $2$-form on $M$, and $\tilde{\lambda}$ its microlocal lift to $S^{*}M$. All contractions and norms are taken with respect to the effective metric $g^{\prime}$.
@@
Defining the fundamental local geometric quantity governing the analysis, the microlocal amplitude of $\tilde{\lambda}$ is defined pointwise by the contraction
@@
where $\tilde{\lambda}^{\sharp}$ denotes the metric dual with respect to $g^{\prime}$. This quantity measures the intensity of the lifted differential form along a given cotangent direction. Unlike classical norms, it retains full directional information encoded in the cosphere bundle. The corresponding global geometric functional is obtained by pairing this density with the canonical Liouville volume form on $S^{*}M$,
@@ V.1 Energy Functionals
where $X_{u}$ denotes the canonical lifted transport vector field on $S^{*}M$, $\tilde{R}$ is a zeroth-order curvature operator induced by the effective connection, $\tilde{C}$ collects commutator terms arising from microlocalization, and $\tilde{F}$ represents external forcing. Doing a energy functional approach to this equation to have boundary conditions, we define the lifted quadratic energy functional as
$\mathcal{E}(t):=\frac{1}{2}\int_{S^{*}M}\langle\tilde{\lambda}(t),\tilde{\lambda}(t)\rangle_{g^{\prime}}\,\mu_{S^{*}M}$ (92)
Differentiating in time along solutions of (91) yields
$\frac{d}{dt}\mathcal{E}(t)$ $=\int_{S^{*}M}\langle\partial_{t}\tilde{\lambda},\tilde{\lambda}\rangle_{g^{\prime}}\,\mu_{S^{*}M}+\frac{1}{2}\int_{S^{*}M}\partial_{t}\langle\tilde{\lambda},\tilde{\lambda}\rangle_{g^{\prime}}\,\mu_{S^{*}M}$ (93)
If the effective metric $g^{\prime}$ is fixed or depends smoothly on $u$, the second term is controlled by $\|\nabla u\|_{L^{\infty}}$ and can be absorbed into the estimates below.
Using (91), we estimate each contribution separately. The transport term satisfies
$\left\|\int_{S^{*}M}\langle\mathcal{L}_{X_{u}}\tilde{\lambda},\tilde{\lambda}\rangle_{g^{\prime}}\,\mu_{S^{*}M}\right\|\leq\sigma\|\nabla u\|_{L^{\infty}}\|\tilde{\lambda}\|_{L^{2}(S^{*}M)}^{2}$ (94)
while the dissipative term is nonnegative,
$\int_{S^{*}M}\langle\nu\|\xi\|_{g^{\prime}}^{2}\tilde{\lambda},\tilde{\lambda}\rangle_{g^{\prime}}\,\mu_{S^{*}M}=\nu\int_{S^{*}M}\|\xi\|_{g^{\prime}}^{2}\|\tilde{\lambda}\|_{g^{\prime}}^{2}\,\mu_{S^{*}M}\geq 0$ (95)
The curvature term admits the pointwise bound
$\|\langle\tilde{R}[\tilde{\lambda}],\tilde{\lambda}\rangle_{g^{\prime}}\|\leq\|\operatorname{Ric}^{u}\|_{L^{\infty}(S^{*}M)}\|\tilde{\lambda}\|_{g^{\prime}}^{2}$ (96)
The commutator term, arising from non-exact microlocalization, satisfies
$\|\langle\tilde{C}[\tilde{\lambda}],\tilde{\lambda}\rangle\|\leq\sigma\,\|\nabla u\|_{L^{\infty}(M)}\,\|\tilde{\lambda}\|_{g^{\prime}}^{2}$ (97)
Finally, the forcing term is controlled linearly,
@@
Collecting all contributions, we obtain
$\frac{d}{dt}\mathcal{E}(t)\leq\sigma\bigl(1+\|\nabla u\|_{L^{\infty}}+\|\nabla^{2}u\|_{L^{\infty}}\bigr)\mathcal{E}(t)$ (99)
where $\sigma>0$ depends only in the estimates depend on $\|\nabla u\|_{L^{\infty}}$ and the background curvature, but remain bounded on finite time intervals under the regularity assumptions considered. By Grönwall’s inequality,
$\mathcal{E}(t)\leq\mathcal{E}(t_{0})\exp\!\left(\sigma\int_{t_{0}}^{t}\bigl(1+\|\nabla u(s)\|_{L^{\infty}}\\ +\|\nabla^{2}u(s)\|_{L^{\infty}}\bigr)\,ds\right)$ (100)
In particular, as long as the velocity field $u$ remains bounded in $W^{2,\infty}$, the lifted energy $\mathcal{E}(t)$ cannot blow up in finite time. Moreover, the dissipative term provides additional control of high–frequency microlocal components. To quantify those directional concentration effects beyond quadratic energy, we introduce the microlocal functional
$\mathcal{V}_{\lambda}(t):=\int_{S^{*}M}\|\tilde{\lambda}(x,\xi,t)\|_{g^{\prime}}\,\mu_{S^{*}M}$ (101)
Since the Liouville measure is invariant under the flow generated by $X_{u}$,
@@
integration of the pointwise balance law associated with (91) yields
$\frac{d}{dt}\mathcal{V}_{\lambda}(t)+\nu\int_{S^{*}M}\|\xi\|_{g^{\prime}}^{2}\|\tilde{\lambda}\|_{g^{\prime}}^{2}\,\mu_{S^{*}M}\\ \leq\sigma\|\nabla u\|_{L^{\infty}}\mathcal{V}_{\lambda}(t)+\|\tilde{F}\|_{L^{1}(S^{*}M)}$ (103)
This inequality expresses a purely geometric fact, the microlocal energy cannot accumulate along cotangent directions without being dissipated, unless the symmetric deformation of the flow becomes dominant. Decomposing $\nabla u=S+\Omega$ into symmetric and antisymmetric parts, one obtains along integral curves of $X_{u}$ the pointwise inequality
$\frac{d}{dt}\|\tilde{\lambda}\|\leq\|S(t)\|_{L^{\infty}}\|\tilde{\lambda}\|-\nu\|\tilde{\lambda}\|$ (104)
with solution
$\|\tilde{\lambda}(t)\|\leq\|\tilde{\lambda}(t_{0})\|\exp\!\left[\int_{t_{0}}^{t}(\|S(s)\|_{L^{\infty}}-\nu)\,ds\right]$ (105)

### v3

@@ IV.5 Solution Formula and Reconstruction
  • $\mathcal{T}$ is the time-ordering operator,
  • $G(t,s)$ is the Green operator constructed above,
  • $\tilde{\phi}_{0}$ is the initial microlocal data,
  • $\tilde{R}$ and $\tilde{F}$ are the curvature and forcing terms.
@@
Finally, the physical solution is recovered by integrating over the cotangent directions:
$\phi(x,t)=\int_{S^{*}_{x}M}\tilde{\phi}(x,t,\xi)\,d\xi$ (100)
Proposition 4 (Well-definedness of reconstruction).
If $\tilde{\phi}\in L^{\infty}([t_{0},t];L^{2}(S^{*}M))\cap L^{\infty}(S^{*}M)$, then the reconstruction formula (100) is well-defined for almost every $x\in M$, and the reconstructed field $\phi$ belongs to $L^{\infty}([t_{0},t];L^{2}(M))$.
Proof.
Since $S^{*}_{x}M$ is compact and $\tilde{\phi}$ is $L^{2}$-integrable on fibers, the integral in (100) is well-defined for each $x$ by the Cauchy-Schwarz inequality. The $L^{2}(M)$ norm of $\phi$ is controlled by integrating the fiber norms over $M$, which is bounded by the total $L^{2}(S^{*}M)$ norm of $\tilde{\phi}$ via Fubini’s theorem. ∎
@@
The reconstruction equation (100) recovers $\phi$ in a weak or averaged sense over directions, which is sufficient for the analysis of regularity.
IV.6 Compatibility with Vorticity Evolution
Analogously, we apply the same procedure to the vorticity 2-form, obtaining the microlocal vorticity equation:
$\partial_{t}\tilde{\omega}+\mathcal{L}_{X_{u}}\tilde{\omega}=-\nu\|\xi\|^{2}_{g^{\prime}}\tilde{\omega}+\tilde{R}_{\omega}+i\xi\wedge\tilde{F},\qquad\xi\wedge\tilde{\omega}=0$ (101)
@@ IV.6 Compatibility with Vorticity Evolution
The condition $\xi\wedge\tilde{\omega}=0$ is the microlocal version of the closure condition $d\omega=0$. Equation (101) is linear and dissipative, and its solution is formally written as
$\tilde{\omega}(t)=\mathcal{T}\left\{\exp\left[-\int_{t_{0}}^{t}\left(\mathcal{L}_{X_{u}(s)}+\nu\|\xi\|^{2}_{g^{\prime}}\right)ds\right]\right\}\tilde{\omega}_{0}\\ +\int_{t_{0}}^{t}G(t,s)\left[\tilde{R}_{\omega}(s)+i\xi\wedge\tilde{F}(s)\right]ds$ (102)
Proposition 5 (Compatibility of velocity and vorticity lifts).
If $\tilde{\phi}(t)$ is a solution to (75), then $\tilde{\omega}(t):=d\tilde{\phi}(t)$ satisfies the microlocal vorticity equation (101).
Proof.
This follows from the commutator estimates established in Proposition 2. Applying the exterior derivative $d$ to (75) and using $d^{2}P=0$ yields
$\partial_{t}(d\tilde{\phi})+[d,\mathcal{L}_{X_{u}}]\tilde{\phi}+\mathcal{L}_{X_{u}}(d\tilde{\phi})+[d,\nu\|\xi\|^{2}_{g^{\prime}}]\tilde{\phi}+\nu\|\xi\|^{2}_{g^{\prime}}(d\tilde{\phi})\\ =d\tilde{R}+d\tilde{F}$ (103)
Using the identity $d,\mathcal{L}_{X_{u}}]=\mathcal{L}_{dX_{u}}$ and absorbing the commutator terms into the residual $\tilde{R}_{\omega}$, we recover (101). ∎
The reconstructed vorticity is given by
$\omega(x,t)=\int_{S^{*}_{x}M}\tilde{\omega}(x,t,\xi)\,d\xi$ (104)
@@
This formulation reduces the global regularity problem to the control and convergence of geometric microlocal invariants, which is the subject of the following section.
V Stability, Energy, and Geometric Criteria for Regularity and Smoothness
@@
In the previous sections, the incompressible Navier–Stokes equations were reformulated as a linear, non-autonomous, dissipative dynamical system acting on microlocal lifts of differential forms over the compact phase space $S^{*}M$. We now exploit this structure to derive quantitative stability estimates and geometric criteria for regularity. The central idea is that any potential singular behavior of the physical solution must manifest itself as a geometric instability of the microlocal dynamics on $S^{*}M$. Viscosity, compactness, and the geometric nature of the lift combine to impose strong restrictions on such instabilities.
@@
Throughout this section, $(M,g)$ denotes a smooth Riemannian manifold and $S^{*}M$ its unit cosphere bundle endowed with the canonical Liouville measure $\mu_{S^{*}M}$. We assume that the velocity field satisfies
$u\in L^{\infty}([t_{0},t);W^{2,\infty}(M))$ (105)
@@
and that all geometric quantities associated with the effective connection $\nabla^{u}$ remain smooth on $t_{0},t)\times S^{*}M$.
Let $\lambda\in\{\phi,\omega\}$ denote either the velocity 1-form or the vorticity 2-form on $M$, and $\tilde{\lambda}$ its microlocal lift to $S^{*}M$. All contractions and norms are taken with respect to the effective metric $g^{\prime}$.
V.1 Microlocal Amplitude and Energy Functionals
@@
The fundamental local geometric quantity governing the analysis is the microlocal amplitude of $\tilde{\lambda}$, defined pointwise by the contraction
@@
where $\tilde{\lambda}^{\sharp}$ denotes the metric dual with respect to $g^{\prime}$. This quantity measures the intensity of the lifted differential form along a given cotangent direction $\xi$. Unlike classical norms on $M$, it retains full directional information encoded in the cosphere bundle.
The corresponding global geometric functional is obtained by pairing this density with the canonical Liouville volume form on $S^{*}M$:
@@
$\partial_{t}\tilde{\lambda}+\mathcal{L}_{X_{u}}\tilde{\lambda}+\nu\|\xi\|^{2}_{g^{\prime}}\tilde{\lambda}=\tilde{R}[\tilde{\lambda}]+\tilde{C}[\tilde{\lambda}]+\tilde{F}$ (108)
where $X_{u}$ denotes the canonical lifted transport vector field on $S^{*}M$, $\tilde{R}$ is a zeroth-order curvature operator induced by the effective connection, $\tilde{C}$ collects commutator terms arising from microlocalization, and $\tilde{F}$ represents external forcing.
We define the lifted quadratic energy functional as
$E(t):=\frac{1}{2}\int_{S^{*}M}\langle\tilde{\lambda}(t),\tilde{\lambda}(t)\rangle_{g^{\prime}}\,\mu_{S^{*}M}$ (109)
Theorem 1 (Geometric energy inequality).
Let $(M,g)$ be a smooth Riemannian manifold and let $u\in L^{\infty}([t_{0},t];W^{2,\infty}(M))$ be divergence-free. Let $\tilde{\lambda}$ solve the lifted evolution equation (108) and define $E(t)$ as in (109). Then there exists $\sigma>0$, depending only on the background geometry, such that
$E(t)\leq E(t_{0})\exp\left[\sigma\int_{t_{0}}^{t}\left(1+\|\nabla u(s)\|_{L^{\infty}}+\|\nabla^{2}u(s)\|_{L^{\infty}}\right)ds\right]$ (110)
Proof.
Differentiating $E(t)$ in time along solutions of (108):
$\frac{d}{dt}E(t)=\int_{S^{*}M}\langle\partial_{t}\tilde{\lambda},\tilde{\lambda}\rangle_{g^{\prime}}\,\mu_{S^{*}M}+\frac{1}{2}\int_{S^{*}M}\partial_{t}\langle\tilde{\lambda},\tilde{\lambda}\rangle_{g^{\prime}}\,\mu_{S^{*}M}$ (111)
If the effective metric $g^{\prime}$ is fixed or depends smoothly on $u$, the second term is controlled by $\|\nabla u\|_{L^{\infty}}$ and can be absorbed into the estimates below. More precisely, since $g^{\prime}$ depends on the integral of the deformation tensor $S$, we have
$\left|\partial_{t}g^{\prime}_{ik}\right|\leq\alpha\|S(t)\|_{L^{\infty}}\leq\alpha\|\nabla u(t)\|_{L^{\infty}}$ (112)
Using (108), we estimate each contribution separately.
Transport term:
The Lie derivative $\mathcal{L}_{X_{u}}$ along the Hamiltonian vector field $X_{u}$ preserves the Liouville measure $\mu_{S^{*}M}$ up to lower-order corrections. Integration by parts on the compact manifold $S^{*}M$ yields
$\left|\int_{S^{*}M}\langle\mathcal{L}_{X_{u}}\tilde{\lambda},\tilde{\lambda}\rangle_{g^{\prime}}\,\mu_{S^{*}M}\right|\leq\sigma\|\nabla u\|_{L^{\infty}}\|\tilde{\lambda}\|_{L^{2}(S^{*}M)}^{2}$ (113)
where $\sigma$ depends on the geometry of $(M,g)$.
Dissipative term:
The viscous contribution is non-negative:
$\int_{S^{*}M}\langle\nu\|\xi\|^{2}_{g^{\prime}}\tilde{\lambda},\tilde{\lambda}\rangle_{g^{\prime}}\,\mu_{S^{*}M}=\nu\int_{S^{*}M}\|\xi\|^{2}_{g^{\prime}}\|\tilde{\lambda}\|^{2}_{g^{\prime}}\,\mu_{S^{*}M}\geq 0$ (114)
Curvature term:
The operator $\tilde{R}$ is zeroth-order and depends on the effective Ricci curvature. By the controlled geometry theorem established in Section III, the effective Ricci tensor satisfies
$\|\text{Ric}^{u}\|_{L^{\infty}(S^{*}M)}\leq\sigma(1+\|\nabla u\|_{L^{\infty}}+\|\nabla^{2}u\|_{L^{\infty}})$ (115)
Thus the curvature term admits the pointwise bound
$\|\langle\tilde{R}[\tilde{\lambda}],\tilde{\lambda}\rangle_{g^{\prime}}\|\leq\|\text{Ric}^{u}\|_{L^{\infty}(S^{*}M)}\|\tilde{\lambda}\|^{2}_{g^{\prime}}$ (116)
Commutator term:
By Lemma 1, the commutator $\tilde{C}$ satisfies
$\|\langle\tilde{C}[\tilde{\lambda}],\tilde{\lambda}\rangle\|\leq\sigma\|\nabla u\|_{L^{\infty}}\|\tilde{\lambda}\|^{2}_{g^{\prime}}$ (117)
Forcing term:
The external forcing is controlled linearly:
@@
Collecting all contributions and using the Cauchy-Schwarz inequality for the forcing term:
$\frac{d}{dt}E(t)\leq\sigma\left(1+\|\nabla u\|_{L^{\infty}}+\|\nabla^{2}u\|_{L^{\infty}}\right)E(t)\\ +\frac{1}{2\nu}\|\tilde{F}\|_{L^{2}(S^{*}M)}^{2}$ (119)
If the forcing is regular (or zero), the second term is controlled independently. Applying Grönwall’s inequality yields (110). ∎
Corollary 1 (Geometric regularity criterion).
As long as the velocity field $u$ remains bounded in $W^{2,\infty}(M)$, the lifted energy $E(t)$ cannot blow up in finite time.
V.2 Microlocal Volume Functional and Directional Control
To quantify directional concentration effects beyond quadratic energy, we introduce the microlocal volume functional:
$V_{\lambda}(t):=\int_{S^{*}M}\|\tilde{\lambda}(x,\xi,t)\|_{g^{\prime}}\,\mu_{S^{*}M}$ (120)
This functional measures the total $L^{1}$ norm of the microlocal amplitude over the phase space.
Proposition 6 (Dissipation inequality for the volume functional).
The functional $V_{\lambda}(t)$ satisfies the dissipation inequality
$\frac{d}{dt}V_{\lambda}(t)+\nu\int_{S^{*}M}\|\xi\|^{2}_{g^{\prime}}\|\tilde{\lambda}\|^{2}_{g^{\prime}}\,\mu_{S^{*}M}\leq\sigma\|\nabla u\|_{L^{\infty}}V_{\lambda}(t)\\ +\|\tilde{F}\|_{L^{1}(S^{*}M)}$ (121)
Proof.
Since the Liouville measure is invariant under the flow generated by $X_{u}$:
@@ V.2 Microlocal Volume Functional and Directional Control
Integration of the pointwise balance law associated with (108) yields:
$\frac{d}{dt}V_{\lambda}(t)$ $=\int_{S^{*}M}\frac{d}{dt}\|\tilde{\lambda}\|_{g^{\prime}}\,\mu_{S^{*}M}$ (123)
$=\int_{S^{*}M}\frac{\langle\partial_{t}\tilde{\lambda},\tilde{\lambda}\rangle_{g^{\prime}}}{\|\tilde{\lambda}\|_{g^{\prime}}}\,\mu_{S^{*}M}$ (124)
Substituting the evolution equation and using the divergence theorem on the compact manifold $S^{*}M$ (which has no boundary), the transport term $\mathcal{L}_{X_{u}}$ integrates to zero up to curvature corrections. The dissipative term is negative:
$\int_{S^{*}M}\frac{\nu\|\xi\|^{2}_{g^{\prime}}\|\tilde{\lambda}\|^{2}_{g^{\prime}}}{\|\tilde{\lambda}\|_{g^{\prime}}}\,\mu_{S^{*}M}=\nu\int_{S^{*}M}\|\xi\|^{2}_{g^{\prime}}\|\tilde{\lambda}\|_{g^{\prime}}\,\mu_{S^{*}M}$ (125)
The curvature and commutator terms contribute lower-order corrections bounded by $\sigma\|\nabla u\|_{L^{\infty}}V_{\lambda}(t)$. ∎
This inequality expresses a purely geometric fact: the microlocal energy cannot accumulate along cotangent directions without being dissipated, unless the symmetric deformation of the flow becomes dominant.
V.3 Pointwise Stretching Estimate
Decomposing the velocity gradient as $\nabla u=S+\Omega$ into symmetric and antisymmetric parts, we obtain along integral curves of $X_{u}$ the pointwise differential inequality:
$\frac{d}{dt}\|\tilde{\lambda}\|_{g^{\prime}}\leq\|S(t)\|_{L^{\infty}}\|\tilde{\lambda}\|_{g^{\prime}}-\nu\|\tilde{\lambda}\|_{g^{\prime}}$ (126)
Proof.
Along a microlocal trajectory $\gamma(\tau)=\Psi_{\tau,t_{0}}(x_{0},\xi_{0})$, we have
$\frac{d}{d\tau}\|\tilde{\lambda}(\gamma(\tau))\|_{g^{\prime}}=\langle\mathcal{L}_{X_{u}}\tilde{\lambda},\tilde{\lambda}\rangle_{g^{\prime}}/\|\tilde{\lambda}\|_{g^{\prime}}$ (127)
The Lie derivative decomposes as